OpenStudy (jackellyn):

Solve the DE using Laplace Transforms y''+9y=0 y(0)=0 y'(0)=2

OpenStudy (jackellyn):

So far I have gotten to this: \[L(y)=2 \div (s^2+9)\]

OpenStudy (blurbendy):

do you have to use Laplace transforms

zepdrix (zepdrix):

Hmm ok your work so far looks good. Let's multiply the top and bottom by 3,\[\large \mathscr{L}\left[y\right]=\frac{3}{3}\cdot\frac{2}{s^2+9}\]Which will give us,\[\large \mathscr{L}\left[y\right]=\frac{2}{3}\cdot\color{green}{\frac{3}{s^2+3^2}}\] Now you're able to take the inverse laplace transform.

zepdrix (zepdrix):

That green term should look like a familiar trig laplace transform

OpenStudy (jackellyn):

@zepdrix Thanks, but what about the initial conditions? I don't really see any constants to solve for.

OpenStudy (jackellyn):

I got y(t)=2/3 sin (3t)